# Graph Laplacian-based Bayesian Multi-fidelity Modeling

### v1

@@ 5.4 Composite cantilever beam
This problem is described in Cheng et al. [30] and the authors have also provided the data. It involves a composite cantilever beam subject to uniform distributed vertical load. As shown in Figure 10, the cross section of the beam is composed of three materials with different properties, and there are five holes running through the lateral extent of the beam.
Figure 10: Schematic diagram of the composite cantilever beam subject to a uniform vertical load (Case 4).
@@
The low-fidelity model is an analytical expression derived from the Euler–Bernoulli beam theory, which ignores the out of plane deformation of the beam and the effect of the holes in the beam. The high-fidelity model is based on the solution of the equations of three-dimensional elasticity using the finite element method. The reader is referred to [30] for details of the problem and solution techniques.
@@
The low-fidelity dataset consists of $N=10,000$ data points, and the high-fidelity points used to compute the multi-fidelity estimates contains $M=50$ (0.5% of the total) data points. We do not normalize the data in this case. In Figure 11 we plot the low-, high-, and multi-fidelity versions of vertical displacements for four instances of input parameters. In all cases, the low-fidelity model under-predicts the displacement, while the multi-fidelity approach is able to correct this. Further, we observe that the multi-fidelity approach is able to capture the effect of the circular holes in the beam, represented by undulations in the displacement, which is missing from the low-fidelity displacement.
In Figure 12 we have plotted the histogram for the error distributions (defined in Eq. (5.2)) for the low-and multi-fidelity data. The improvement in the performance of the multi-fidelity approach is significant, with no overlap between the low- and multi-fidelity errors. The mean values of the low- and multi-fidelity errors are 30.7% and 4.43%, respectively (see Table 3), which implies a 85.6% reduction in error due to the multi-fidelity approach.
Figure 11: Instances of low-, multi-, and high-fidelity solutions for the composite beam problem (Case 4). We note that in some cases the multi- and high-fidelity curves almost completely overlap.
Figure 12: Error distribution in the low- and multi-fidelity data for the composite beam problem (Case 4).
@@ 5.5 Heat flux in cavity flow
Our final example is also solved using data from Cheng et al. [30] and involves predicting the distribution of heat flux in a fluid. In this problem the domain is a closed cavity in the form of a unit square which contains a fluid. The left wall of the cavity is maintained at a fixed temperature $T_{h}$, while on the the right wall (the cold wall) a stochastic profile, $T_{c}(y)$, with mean value $\bar{T}_{c}<T_{h}$ is prescribed. The two horizontal walls are assumed to be adiabatic. The temperature difference between the vertical walls onsets a clockwise motion of the fluid inside the cavity, which is modeled using the unsteady, incompressible Navier-Stokes equations coupled with an equation for the conservation of energy. No-slip boundary conditions are prescribed for the velocity on the walls of the cavity.
Figure 13: Schematic of temperature-driven cavity flow (Case 5).
@@
The low- and high-fidelity data are obtained by solving the Navier-Stokes equations using a finite volume method with $16\times 16$ and $256\times 256$ grids, respectively. The reader is referred to [30] for more details regarding the problem and the data generation process.
@@
The low-fidelity dataset contains $N=10,000$ data points, and $M=100$ high-fidelity points are utilized to determine the multi-fidelity model (1% of the total). We note that each data point is obtained by solving the Navier-Stokes equations. Also in this case, we do not employ any data normalization procedure. In Figure 14, we have plotted four instances of the low-, high and multi-fidelity heat flux distributions. In each case we note that the low-fidelity solution over-predicts the heat flux and is unable to capture the subtle variations as a function of the vertical coordinate. The multi-fidelity approach improves on both these aspects and is much closer to the reference high-fidelity solution.
In Figure 15 we plot the histogram for the error distributions (defined in Eq. (5.2)) for the low-and multi-fidelity data for this problem. Once again, the multi-fidelity distribution is concentrated in a region where the error is much smaller. The mean values of the low- and multi-fidelity errors are 18.63% and 3.87%, respectively (see Table 3), which implies a percentage reduction in error of 79.2%.
Figure 14: Instances of low-, multi-, and high-fidelity solutions for the cavity flow problem (Case 5).
Figure 15: Distribution of error in the low- and high-fidelity data for cavity flow problem (Case 5).
@@ 5.6 Uncertainty of the multi-fidelity estimates
In this section we visualize the uncertainty distribution of the multi-fidelity estimates for all considered problems. Specifically, we utilize the Uniform Manifold Approximation and Projection (UMAP) method [65] to project the multi-fidelity data onto a two-dimensional plane. UMAP is a dimensionality reduction technique that transforms high-dimensional data into a lower-dimensional space, preserving the local structure by keeping similar data points close together and dissimilar ones farther apart.
@@
Figure 16: UMAP embedding of multi-fidelity data (Case 1 to 5) into a plane. Data points are colored based on the variance determined by the multi-fidelity method.
@@ 6 Conclusions
In this study, we propose a Bayesian extension of the graph Laplacian-based spectral multi-fidelity model (SpecMF) [42]. By leveraging the spectral properties of the graph Laplacian, this method constructs a prior distribution that captures the underlying structure of the data, which is then refined using a likelihood function determined using a few select high-fidelity data points. It is shown that resulting posterior distribution for the multi-fidelity coordinates of the data points is a multivariate Gaussian distribution, and the mean and the covariance of this distribution can be determined by solving linear systems of equations. We present two efficient numerical methods for solving these systems. We apply the method to a wide range of problems in solid and fluid mechanics where the quantities of interest range from low-dimensional vectors ($D=\mathcal{O}(10)$) to discrete representations of two-dimensional fields ($D=\mathcal{O}(10^{4})$). In all cases the multi-fidelity approach improves the accuracy of the underlying low-fidelity data by 75 to 85%, while only using a small proportion of high-fidelity data (from 0.5 to 3.3%).
@@ Appendix A General graph Laplacian normalization
In the above, we have assumed that $p=q$ in the definition of $\bm{L}$ in (2.3). Other normalizations of the graph Laplacian are possible and indeed common for different application settings, for example $(p,q)=(1,0)$ gives the random walk Laplacian, so-called because $\bm{D}^{-1}\bm{W}$ is the transition matrix of a random walk on the graph. It is well-known that the choice of normalization has a direct impact on the output of any graph Laplacian based algorithm, and a spectral analysis of the effects of the choice of normalization in the large-data limit is given in Hoffmann et al. [46].
@@
$\langle\bm{u},\bm{v}\rangle_{p-q}$ $:=\bm{u}^{T}\bm{D}^{p-q}\bm{v}$ $\bm{u},\bm{v}\in\mathbb{R}^{N},$
$\langle\bm{A},\bm{B}\rangle_{p-q,F}$ $:=\operatorname{tr}\left(\bm{A}^{T}\bm{D}^{p-q}\bm{B}\right)$ $\bm{A},\bm{B}\in\mathbb{R}^{N\times K},$
@@
We then redefine our prior (2.4) for $\bm{\Phi}$ using the reweighted Frobenius norm, i.e.,
$p(\bm{\Phi})\propto\exp\left(-\frac{\omega}{2}\langle\bm{\Phi},\left(\bm{L}^{(p,q)}+\tau\bm{I}_{N}\right)^{\beta}\bm{\Phi}\rangle_{p-q,F}\right).$
Since the inner product (by design) respects the orthonormality of the eigenvectors, the expression for the prior in terms of the coefficient matrix (2.5) is unchanged. Choosing the likelihood as in (2.6), we can express the posterior as in (2.8), except now
$\bm{C}$ $=\left(\frac{1}{\sigma^{2}}\bm{P}_{M}^{T}\bm{P}_{M}+\omega\bm{D}^{p-q}(\bm{L}^{(p,q)}+\tau\bm{I}_{N})^{\beta}\right)^{-1},$ (A.1)
@@
$\left(\frac{1}{\sigma^{2}}\bm{P}_{M}^{T}\bm{P}_{M}+\omega{\bm{D}^{p-q}}(\bm{L}^{(p,q)}+\tau\bm{I}_{N})^{\beta}\right)\bm{\Phi}^{*}=\frac{1}{\sigma^{2}}\bm{P}_{M}^{T}\hat{\bm{\Phi}}.$ (A.2)
@@
$a\bm{I}_{N}-\bm{L}^{((p+q)/2,(p+q)/2)},$
before being converted to low-lying $\langle\cdot,\cdot\rangle_{p-q}$-orthonormal eigenvectors of $\bm{L}^{(p,q)}$ by multiplying each leading eigenvector of $a\bm{I}_{N}-\bm{L}^{((p+q)/2,(p+q)/2)}$ by $\bm{D}^{-(p-q)/2}$. This $\langle\cdot,\cdot\rangle_{p-q}$-orthonormality, combined with the redefinition of the prior density above, ensures that the eventual expressions for $\bm{C}_{A_{K}}$ and $\bm{A}^{*}_{K}$ in (3.2) and (3.3) are unchanged, except that $\bm{\Psi}_{K}$ now refers to the low-lying eigenvectors of $\bm{L}^{(p,q)}$.
@@
$\tilde{\bm{U}}\bm{\Sigma}\tilde{\bm{U}}^{T}=\hat{\bm{D}}^{-\frac{1}{2}}\bm{W}(:,X)\bm{W}(X,X)^{\dagger}\bm{W}(X,:)\hat{\bm{D}}^{-\frac{1}{2}},$
@@
$\bm{L}^{(p,1-p)}\approx\bm{I}_{N}-\hat{\bm{D}}^{\frac{1}{2}-p}\tilde{\bm{U}}\bm{\Sigma}\tilde{\bm{U}}^{T}\hat{\bm{D}}^{p-\frac{1}{2}}=\bm{I}_{N}-\bm{U}\bm{\Sigma}\bm{V}^{T},$
where $\bm{U}:=\hat{\bm{D}}^{\frac{1}{2}-p}\tilde{\bm{U}}$, $\bm{V}:=\hat{\bm{D}}^{p-\frac{1}{2}}\tilde{\bm{U}}$, and $\bm{V}^{T}\bm{U}=\bm{I}_{K}$. The approximation of $(\bm{L}^{(p,1-p)}+\tau\bm{I}_{N})^{\beta}$ then goes through with $\tilde{\bm{U}}$ replaced by $\bm{U}$ and $\tilde{\bm{U}}^{T}$ replaced by $\bm{V}^{T}$. Redefining (in line with the $\bm{D}^{p-q}$ term in (A.1))
$\bm{\Theta}:=\bm{P}_{M}^{T}\bm{P}_{M}+\sigma^{2}\omega(1+\tau)^{\beta}\hat{\bm{D}}^{2p-1}\in\mathbb{R}^{N\times N},$
it follows that Equation A.2 can be approximately solved by solving
$\left(\bm{\Theta}-\hat{\bm{D}}^{2p-1}\bm{U\Xi V}^{T}\right)\bm{\Phi}^{*}=\left(\bm{\Theta}-\bm{V\Xi V}^{T}\right)\bm{\Phi}^{*}=\bm{P}_{M}^{T}\hat{\bm{\Phi}},$ (A.3)
@@
$\bm{C}=\left(\frac{1}{\sigma^{2}}\bm{P}_{M}^{T}\bm{P}_{M}+\omega\bm{D}^{2p-1}(\bm{L}+\tau\bm{I}_{N})^{\beta}\right)^{-1}\approx\sigma^{2}\left(\bm{\Theta}-\bm{V}\bm{\Xi V}^{T}\right)^{-1}.$
Hence, the methods for solving (A.3) and computing $\bm{C}$ are as before, except for $\tilde{\bm{U}}$ everywhere replaced by $\bm{V}$. The computational complexities in Section 3 are unaffected by any of these changes.
Finally, in Section 4, to ensure that the minimizer of the objective function solves (A.2), the regularizer must be redefined as
$\mathcal{R}(\bm{\Theta}):=\langle\mathbf{\Theta},(\bm{L}^{(p,q)}+\tau\bm{I}_{N})^{\beta}\mathbf{\Theta}\rangle_{p-q,F},$

### v2

@@ 5.4 Composite cantilever beam
This problem is described in Cheng et al. [CHENG2024116793] and the authors have also provided the data. It involves a composite cantilever beam subject to uniform distributed vertical load. As shown in Figure 10, the cross section of the beam is composed of three materials with different properties, and there are five holes running through the lateral extent of the beam.
Figure 10: Schematic diagram of the composite cantilever beam subject to a uniform vertical load (Case 4). The presence of the holes is neglected in the low-fidelity model, where an analytical expression for the displacement is derived based on the Euler-Bernoulli beam theory. The parameters include the Young’s modulus of the three materials composing the beam, and the vertical load applied.
@@
The low-fidelity model is an analytical expression derived from the Euler–Bernoulli beam theory, which ignores the out of plane deformation of the beam and the effect of the holes in the beam. The high-fidelity model is based on the solution of the equations of three-dimensional elasticity using the finite element method. Because the low-fidelity model is an analytical model, its computational cost is negligible. The reader is referred to [CHENG2024116793] for details of the problem and solution techniques.
@@
The low-fidelity dataset consists of $N=$ 10,00010{,}000 data points, and the high-fidelity points used to compute the multi-fidelity estimates contains $M=$ $50$ ($0.5\%$ of the total) data points. We do not normalize the data in this case. In Figure 11 we plot the low-, high-, and multi-fidelity versions of vertical displacements for four instances of input parameters. In all cases, the low-fidelity model under-predicts the displacement, while the multi-fidelity approach is able to correct this. Further, we observe that the multi-fidelity approach is able to capture the effect of the circular holes in the beam, represented by undulations in the displacement, which is missing from the low-fidelity displacement.
[…]
Figure 11: Instances of low-, multi-, and high-fidelity solutions for the vertical displacement for the composite cantilever beam (Case 4), resulting from different input parameters. The low-fidelity model underestimates the vertical displacement and neglects structural features like holes. On the other hand, the multi-fidelity model provides solutions that closely align with high-fidelity results, sometime almost completely overlapping with the high-fidelity curves.
Figure 12: Histogram of error distribution for low- and multi-fidelity datasets for the composite beam problem (Case 4). The multi-fidelity approach results in a distribution that is narrower and closer to zero. The mean and standard deviation for the low-fidelity distribution are 30.730.7% and $2.68$, respectively. Whereas for the multi-fidelity model the mean and standard deviation are 4.434.43% and $3.34$, respectively.
@@ 5.5 Heat flux in cavity flow
Our final example is also solved using data from Cheng et al. [CHENG2024116793] and involves predicting the distribution of heat flux in a fluid. In this problem the domain is a closed cavity in the form of a unit square which contains a fluid. The left wall of the cavity is maintained at a fixed temperature $T_{h}$, while on the the right wall (the cold wall) a stochastic profile, $T_{c}(y)$, with mean value $\bar{T}_{c}<T_{h}$ is prescribed. The two horizontal walls are assumed to be adiabatic. The temperature difference between the vertical walls onsets a clockwise motion of the fluid inside the cavity, which is modeled using the unsteady, incompressible Navier-Stokes equations coupled with an equation for the conservation of energy. No-slip boundary conditions are prescribed for the velocity on the walls of the cavity.
Figure 13: Schematic of temperature-driven cavity flow (Case 5). The difference of temperature of the vertical walls onsets a clock-wise motion of the fluid inside the cavity. The parameters are the temperature at the hot left wall, and the distribution of temperature on the cold right wall. The quantity of interest is the heat flux through the hot wall at steady state as a function of the vertical coordinate.
@@
The low- and high-fidelity data are obtained by solving the Navier-Stokes equations using a finite volume method with $16\times 16$ and $256\times 256$ grids, respectively. The ratio between the computational costs of the two models is approximately 9,400 [CHENG2024116793]. The reader is referred to [CHENG2024116793] for more details regarding the problem and the data generation process.
@@
The low-fidelity dataset contains $N=$ 10,00010{,}000 data points, and $M=$ $100$ high-fidelity points are utilized to determine the multi-fidelity model ($1\%$ of the total). We note that each data point is obtained by solving the Navier-Stokes equations. Also in this case, we do not employ any data normalization procedure. In Figure 14, we have plotted four instances of the low-, high and multi-fidelity heat flux distributions. In each case we note that the low-fidelity solution over-predicts the heat flux and is unable to capture the subtle variations as a function of the vertical coordinate. The multi-fidelity approach improves on both these aspects and is much closer to the reference high-fidelity solution.
[…]
Figure 14: Instances of heat flux curves corresponding to different input parameters for low-, multi-, and high-fidelity models for the cavity flow problem (Case 5). The multi-fidelity approach results in curves that are closer in magnitude and trend to the high-fidelity solutions.
Figure 15: Histogram of error distribution for low- and multi-fidelity datasets for the cavity flow problem (Case 5). The mean and standard deviation for the low-fidelity distribution are 18.6318.63% and $5.96$, respectively. The mean and standard deviation of the distribution for the multi-fidelity data are 3.873.87% and $2.32$, respectively.
@@ 5.6 Uncertainty of the multi-fidelity estimates
In this section we visualize the uncertainty distribution of the multi-fidelity estimates for all the numerical problems. Specifically, we utilize the Uniform Manifold Approximation and Projection (UMAP) method [McInnes2018] to project the multi-fidelity data onto a two-dimensional plane. UMAP is a dimensionality reduction technique that transforms high-dimensional data into a lower-dimensional space, preserving the local structure by keeping similar data points close together and dissimilar ones farther apart.
@@
Figure 16: The figures show the 2-dimensional UMAP embedding of the multi-fidelity data for the five numerical problems we considered. Here, the nodes of the multi-fidelity graph are mapped into a plane, and the adjacency matrix of the graph is used in the UMAP method to impose a distance metric. This ensures that nodes that are similar are mapped close to each other, and nodes that are dissimilar are mapped far apart. Finally, each data point is colored based on the uncertainty (standard deviation) determined by the multi-fidelity method.
@@ 6 Conclusions
In this study, we propose a Bayesian extension of the graph Laplacian-based spectral multi-fidelity model (SpecMF) [pinti2023graph]. By leveraging the spectral properties of the graph Laplacian, this method constructs a prior distribution that captures the underlying structure of the data, which is then refined using a likelihood function determined using a few select high-fidelity data points. The resulting posterior distribution for the multi-fidelity coordinates of the data points is a multivariate Gaussian distribution, and the mean and the covariance of this distribution can be determined by solving linear systems of equations. We present two efficient numerical methods for solving these systems. We apply the method to a wide range of problems in solid and fluid mechanics where the quantities of interest range from low-dimensional vectors ($D=\mathcal{O}(10)$) to discrete representations of two-dimensional fields ($D=\mathcal{O}(10^{4})$). In all cases the multi-fidelity approach improves the accuracy of the underlying low-fidelity data by 75 to 85%, while only using a small proportion of high-fidelity data (from 0.5 to 3.3%).
We now outline the limitations of the methodology, which can be seen as possible directions for future research. The method depends on the low-fidelity model being able to accurately capture the relationships across different data points leading to an accurate adjacency matrix, which directly affects the multi-fidelity correction. Furthermore, the method may also face challenges when used for hyperbolic problems that exhibit a large Kolmogorov $n$-width, as it may require a substantial number of high-fidelity data points. This is due to the fact that the additive correction used to compute the multi-fidelity estimates is a linear combination of the known low-to-high fidelity displacements, and the number of terms may increasing with increasing Kolmogorov $n$-width. Our approach scales effectively with the dimensionality of the data, which only affects the complexity of computing the edge weights of the graph. [^7] Parallelization strategies for efficiently computing edge weights for very high-dimensional data are discussed in Meng et al. [MengIPOL2017]. Finally, we propose efficient numerical techniques for applying our approach in a large data regime, but future efforts to prove the efficacy of these methods with large datasets should we explored.
@@ Appendix A General graph Laplacian normalization
In the above, we have assumed that $p=q$ in the definition of $\bm{L}$ in (5). Other normalizations of the graph Laplacian are possible and indeed common for different application settings, for example $(p,q)=(1,0)$ gives the random walk Laplacian, so-called because $\bm{D}^{-1}\bm{W}$ is the transition matrix of a random walk on the graph. It is well-known that the choice of normalization has a direct impact on the output of any graph Laplacian based algorithm, and a spectral analysis of the effects of the choice of normalization in the large-data limit is given in Hoffmann et al. [Hoffmann2019SpectralAO].
@@
$\langle\bm{u},\bm{v}\rangle_{p-q}$ $:=\bm{u}^{T}\bm{D}^{p-q}\bm{v}$ $\bm{u},\bm{v}\in\mathbb{R}^{N},$ (72)
$\langle\bm{A},\bm{B}\rangle_{p-q,F}$ $:=\operatorname{tr}\left(\bm{A}^{T}\bm{D}^{p-q}\bm{B}\right)$ $\bm{A},\bm{B}\in\mathbb{R}^{N\times K},$ (73)
@@
We then redefine our prior (8) for $\bm{\Phi}$ using the reweighted Frobenius norm, i.e.,
$p(\bm{\Phi})\propto\exp\left(-\frac{\omega}{2}\langle\bm{\Phi},\left(\bm{L}^{(p,q)}+\tau\bm{I}_{N}\right)^{\beta}\bm{\Phi}\rangle_{p-q,F}\right).$ (76)
Since the inner product (by design) respects the orthonormality of the eigenvectors, the expression for the prior in terms of the coefficient matrix (9) is unchanged. Choosing the likelihood as in (14), we can express the posterior as in (17), except now
$\bm{C}$ $=\left(\frac{1}{\sigma^{2}}\bm{P}_{M}^{T}\bm{P}_{M}+\omega\bm{D}^{p-q}(\bm{L}^{(p,q)}+\tau\bm{I}_{N})^{\beta}\right)^{-1},$ (77)
@@
$\left(\frac{1}{\sigma^{2}}\bm{P}_{M}^{T}\bm{P}_{M}+\omega{\bm{D}^{p-q}}(\bm{L}^{(p,q)}+\tau\bm{I}_{N})^{\beta}\right)\bm{\Phi}^{*}=\frac{1}{\sigma^{2}}\bm{P}_{M}^{T}\hat{\bm{\Phi}}$. (78)
@@
$a\bm{I}_{N}-\bm{L}^{((p+q)/2,(p+q)/2)},$ (79)
before being converted to low-lying $\langle\cdot,\cdot\rangle_{p-q}$-orthonormal eigenvectors of $\bm{L}^{(p,q)}$ by multiplying each leading eigenvector of $a\bm{I}_{N}-\bm{L}^{((p+q)/2,(p+q)/2)}$ by $\bm{D}^{-(p-q)/2}$. This $\langle\cdot,\cdot\rangle_{p-q}$-orthonormality, combined with the redefinition of the prior density above, ensures that the eventual expressions for $\bm{C}_{A_{K}}$ and $\bm{A}^{*}_{K}$ in (30) and (31) are unchanged, except that $\bm{\Psi}_{K}$ now refers to the low-lying eigenvectors of $\bm{L}^{(p,q)}$.
@@
$\tilde{\bm{U}}\bm{\Sigma}\tilde{\bm{U}}^{T}=\hat{\bm{D}}^{-\frac{1}{2}}\bm{W}(:,X)\bm{W}(X,X)^{\dagger}\bm{W}(X,:)\hat{\bm{D}}^{-\frac{1}{2}},$ (80)
@@
$\bm{L}^{(p,1-p)}\approx\bm{I}_{N}-\hat{\bm{D}}^{\frac{1}{2}-p}\tilde{\bm{U}}\bm{\Sigma}\tilde{\bm{U}}^{T}\hat{\bm{D}}^{p-\frac{1}{2}}=\bm{I}_{N}-\bm{U}\bm{\Sigma}\bm{V}^{T},$ (81)
where $\bm{U}:=\hat{\bm{D}}^{\frac{1}{2}-p}\tilde{\bm{U}}$, $\bm{V}:=\hat{\bm{D}}^{p-\frac{1}{2}}\tilde{\bm{U}}$, and $\bm{V}^{T}\bm{U}=\bm{I}_{K}$. The approximation of $(\bm{L}^{(p,1-p)}+\tau\bm{I}_{N})^{\beta}$ then goes through with $\tilde{\bm{U}}$ replaced by $\bm{U}$ and $\tilde{\bm{U}}^{T}$ replaced by $\bm{V}^{T}$. Redefining (in line with the $\bm{D}^{p-q}$ term in (77))
$\bm{\Theta}:=\bm{P}_{M}^{T}\bm{P}_{M}+\sigma^{2}\omega(1+\tau)^{\beta}\hat{\bm{D}}^{2p-1}\in\mathbb{R}^{N\times N},$ (82)
it follows that Equation 78 can be approximately solved by solving
$\left(\bm{\Theta}-\hat{\bm{D}}^{2p-1}\bm{U\Xi V}^{T}\right)\bm{\Phi}^{*}=\left(\bm{\Theta}-\bm{V\Xi V}^{T}\right)\bm{\Phi}^{*}=\bm{P}_{M}^{T}\hat{\bm{\Phi}}$, (83)
@@
$\bm{C}=\left(\frac{1}{\sigma^{2}}\bm{P}_{M}^{T}\bm{P}_{M}+\omega\bm{D}^{2p-1}(\bm{L}+\tau\bm{I}_{N})^{\beta}\right)^{-1}\approx\sigma^{2}\left(\bm{\Theta}-\bm{V}\bm{\Xi V}^{T}\right)^{-1}.$ (84)
Hence, the methods for solving (83) and computing $\bm{C}$ are as before, except for $\tilde{\bm{U}}$ everywhere replaced by $\bm{V}$. The computational complexities in Section 3 are unaffected by any of these changes.
Finally, in Section 4, to ensure that the minimizer of the objective function solves (78), the regularizer must be redefined as
$\mathcal{R}(\bm{\Theta}):=\langle\mathbf{\Theta},(\bm{L}^{(p,q)}+\tau\bm{I}_{N})^{\beta}\mathbf{\Theta}\rangle_{p-q,F},$ (85)
@@
Appendix B Gradient of $\mathcal{J}(\omega)$
We are interested in finding the gradient of function (86) with respect to $\omega$:
$\mathcal{J}(\omega)=\left(\frac{1}{N}\sum_{k=1}^{N}\sqrt{C_{kk}(\omega)}-r\sigma\right)^{2}.$ (86)
Let’s first recall the form of the covariance matrix $\bm{C}$ (18):
$\bm{C}^{-1}=\frac{1}{\sigma^{2}}\bm{P}_{M}^{T}\bm{P}_{M}+\omega(\bm{L}+\tau\bm{I}_{N})^{\beta}.$ (87)
Then, we can write,
$\frac{\mathrm{d}\mathcal{J}}{\mathrm{d}\omega}=\sum_{i,\,j=1}^{N}\frac{\partial\mathcal{J}}{\partial C_{ij}}\frac{\partial C_{ij}}{\partial\omega}.$ (88)
The first term writes
$\frac{\partial\mathcal{J}}{\partial C_{ij}}=\frac{1}{N}\left(\frac{1}{N}\sum_{k=1}^{N}\sqrt{C_{kk}}-r\sigma\right)\frac{1}{\sqrt{C_{ii}}}\delta_{ij},$ (89)
with $\delta_{ij}$ being the Kronecker delta. For the second term, we note that
$\frac{\partial\bm{C}}{\partial\omega}$ $=-\bm{C}\frac{\partial\bm{C}^{-1}}{\partial\omega}\bm{C},$ (90)
$\frac{\partial\bm{C}^{-1}}{\partial\omega}$ $=(\bm{L}+\tau\bm{I}_{N})^{\beta}.$ (91)
(90) is a standard result that can be proven by differentiating both sides of $\bm{C}\bm{C}^{-1}=\bm{I}_{N}$ with respect to $\omega$. Then, we have that
$\frac{\partial C_{ij}}{\partial\omega}=\left[\frac{\partial\bm{C}}{\partial\omega}\right]_{ij}=-\left[\bm{C}(\bm{L}+\tau\bm{I}_{N})^{\beta}\bm{C}\right]_{ij}$ (92)
Hence,
$\frac{\mathrm{d}\mathcal{J}}{\mathrm{d}\omega}=-\left(\frac{1}{N}\sum_{k=1}^{N}\sqrt{C_{kk}}-r\sigma\right)\frac{1}{N}\sum_{i=1}^{N}\frac{1}{\sqrt{C_{ii}}}\left[\bm{C}(\bm{L}+\tau\bm{I}_{N})^{\beta}\bm{C}\right]_{ii}.$ (93)